# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v1

@@ Appendix A Dataset Examples
Figure 1: On the right: the current path for selecting some algorithm(s) based on their validity and performance for a given dataset.On the left: the proposed flip to better understand the space in which ‘good’ datasets exist for an algorithm.
Paragraph 17:
Section 2 describes the structure of the proposed method including its parameters and operators.
Paragraph 20:
In this section, the details of an algorithm that generates data for which a given function, or (equivalently) algorithm, is well-suited is described.This algorithm is to be referred to as “Evolutionary Dataset Optimisation” (EDO).
@@
Figure: Figure 1[Uncaptioned image] Caption: Sample number of rows and columns 2. Sample columns and fill/trim values as needed. Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection. Answer 2: By employing the EDO approach, we can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use. This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy. As a result, the performance metrics of the algorithms improve significantly, leading to more valid outcomes and reliable interpretations. The focus on tailored dataset selection directly contributes to optimal algorithm efficiency, thereby validating the effectiveness of the EDO approach in practical applications.
Appendix B Prompt Engineering
Text-only QA Generation Prompt for the Training Set Task Definition: Generate 3 high quality academic questions and corresponding answers based on the provided text/image/table. Requirements: The questions should require very complex reasoning and global understanding and the answer should be detailed and can be answered by the material, do not add extra information. Remember it is only a paragraph rather than a whole passage so do not ask global question about the study. Generate question-answer pairs instead of multiple-choice questions. Do not generate global question about the paper such as main idea or abstract. You should use English. Expected Output: List your Q and A as: [Q1]: [A1]: [Q2]: [A2]: [Q3]: [A3]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea.
Figure 6: Text-only QA generation prompt for the training set.
Text-image QA Generation Prompt for the Training Set Task Definition: Generate 2 high-quality academic questions and corresponding answers based solely on a provided figure of a research paper, without relying on accompanying text for the questions. However, you may use the provided text to understand the figure. Ensure these questions demand complex reasoning, focusing exclusively on details within the figure that are critical for understanding. The answers must be detailed, directly reflecting the content in the figure and capable of being derived solely from the figure without extraneous information. Requirements: Thoroughly analyze the provided figure of the research paper, attaining a deep understanding of its contents, including but not limited to, the study’s objectives, methodologies, results, and conclusions. Develop 2 academic questions that: Must not mention the name of the figure directly or words like “figure”,“table” in the question. Demand complex reasoning and a comprehensive understanding of the whole figure. Must ask questions that can only be answered by the content in the figure, without reliance on textual information. Integrate knowledge from the entire figure, reflecting a nuanced understanding. Frame questions to elicit detailed and informed responses directly supported by the figure. Craft detailed answers for each question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: List your Q and A as: [Q1]: [Formulate the first academic question here, ensuring it requires complex reasoning and encompasses the entirety of the provided figure for a comprehensive understanding. Based solely on the figure, without referencing accompanying text. Must not mention the name of the figure directly. Must not mention words like “figure”, “table” exc.] [A1]: [Provide a detailed answer here, derived exclusively from the information within the provided figure, ensuring the response is thorough and precise without including extraneous details. Mention from which figure/table you get the answer.] [Q2]: [A2]: Here is the supplementary paragraph text for the figure:
Figure 7: Text-image QA generation prompt for the training set.
Section QA Generation Prompt for the Training Set Task Definition: Generate 3 high-quality academic questions and corresponding answers based on a section of a research paper. Requirements: Each question should generate two types of answers. The first answer should be concise, directly addressing the question with minimal wording. The second answer should include a “chain of thought” that provides a reasoning process and be detailed. You should use English. Generate questions that: Can be answered with simple reasoning and only require a global understanding. The question should be able to answered directly with the material. Do not include 2 or more subquestions in each question. Ensure that the concise answers can be provided using sentences or phrases that do not exceed 20 words in length. Remember it is only a paragraph rather than a whole passage so do not ask global question about the study. Generate question-answer pairs instead of multiple-choice questions. Do not generate global question about the paper such as main idea or abstract. Must not mention the name of the figure directly or words like “figure”, “table” exc in the question. Must ask questions that can only be answered by the content in the figure, without reliance on textual information. Craft the 2 answers for each question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: List your Q and A as: [Q1]: [Insert the academic question here, can be answered directly from the material] [A11]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words] [A12]: [Insert the detailed answer here, based solely on the given information. This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption. Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] [Q2]: [A21]: [A22]: [Q3]: [A31]: [A32]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea. Here is the supplementary paragraph text for the figure:
Figure 8: Section QA generation prompt for the training set.
Cross-paragraph Question Generation Prompt for the Training Set Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question. This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper. Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided). Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper. Note that your question should not directly include the “idx”s of the paragraphs. Expected Output: [Q]: [Your generated question based on the shared themes or information] Here are the selected paragraphs in the paper:
Cross-paragraph Answer Generation Prompt for the Training set Task Definition: Given some selected paragraphs from a research paper, each chosen based on their relevance (excluding the first 3 paragraphs), and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs. The question is crafted to encompass the themes or findings presented in the paragraphs of the chosen paragraphs, aiming for a comprehensive understanding and connection between these elements. Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: [A1]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words.] [A2]: [Insert the detailed answer here, based solely on the given information. This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption. Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] The question is:
@@ Appendix B Prompt Engineering
Figure 9: Cross-paragraph QA generation prompt for the training set.
Text-only Question Generation Prompt for the Test set Task Definition: Create 2 academic questions from a given research paper paragraph. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: No more than 30 words. Incorporate knowledge from the paragraph. Should be answered by text instead of one of the multiple choices. Elicit detailed responses supported by the text. Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
Text-only Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-only Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 10: Text-only type QA generation prompt for the test set.
Text-image Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on the provided figures and tables from a research paper. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: No more than 30 words. Are specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables. Must not mention the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]: [Q2]:
Text-image Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Always use English. Do not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-image Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 11: Text-image QA generation prompt for the test set.
Section Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on a section from a research paper. Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions. You should Always use English. Develop 2 questions that: No more than 30 words. Require an integration of information from all paragraphs and figures/tables in the section. Must not mention the label of the figure/table directly or use words like ‘from the figure/table’. Not based on common knowledge or assumptions not supported by the figures and tables. Expected Output: [Q1]: [Q2]:
Section Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Always use English. Do not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Section Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 12: Section QA generation prompt for the test set.

### v2

@@ Appendix C More Dataset Details
Figure 1: On the right: the current path for selecting some algorithm(s), …
@@
Paragraph 17: Section 2 describes the structure of the proposed method including its parameters and operators. Paragraph 20: In this section, the details of an algorithm that generates data for which a given function, … Figure: Figure 1Refer to caption Caption: Sample number of rows and columns 2. Sample columns and fill/trim values as needed. Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection. Answer 2: We can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use. This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy.
@@
Figure 11: Cross-paragraph Q-E-A triplets generation prompt and data example.
Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches. Note: the -means the results were not reported by that method. Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Appendix D Limitation
For our proposed dataset, current documents are mainly limited to academic papers, so the layout format and subject matter of the documents are relatively simple. We will expand the dataset with more diverse documents. Besides, our model is not specifically designed for some global queries, which will become a key research problem in the future.